# The intersection cohomology Hodge module of toric varieties

### v3

@@ 2.1. Hodge modules
which sits in cohomological degrees $-\dim X,\ldots,0$. Moreover, $F_{\bullet}$ induces a natural filtration on $\operatorname{DR}_{X}(\mathcal{M})$ and the graded pieces $\operatorname{gr}_{k}\operatorname{DR}_{X}\mathcal{M}$ are given by
@@
which also sits in cohomological degrees $-\dim X,\ldots,0$. The maps in this complex are $\mathcal{O}_{X}$-linear and we thus view $\operatorname{gr}_{k}\operatorname{DR}_{X}\mathcal{M}$ as an object in $\mathcal{D}^{b}_{\rm coh}(X)$, the derived category of coherent sheaves of $X$. We mention that even if $X$ is singular, one can define the categories $\operatorname{HM}(X,w)$ and $\operatorname{MHM}(X)$ by embedding into a smooth variety $X\hookrightarrow Y$ and considering the objects in $\operatorname{HM}(Y,w)$ and $\operatorname{MHM}(Y)$, respectively, which are supported on $X$. The categories $\operatorname{HM}(X,w)$ and $\operatorname{MHM}(X)$ do not depend on the choice of the embedding. If $X$ cannot be embedded in a smooth variety, we locally embed each open set and impose suitable compatibility conditions on the intersections. The de Rham complex and the graded pieces $\operatorname{gr}_{k}\operatorname{DR}_{X}\mathcal{M}$ are defined as above by locally embedding $X$ into a smooth variety. The graded pieces of the de Rham complex $\operatorname{gr}_{k}\operatorname{DR}_{X}\mathcal{M}$ also do not depend on the choice of the embedding as objects in $\mathcal{D}_{\operatorname{coh}}^{b}(X)$.
@@ 2.2. Toric varieties
    [Fulton-ToricVar]*Section 3.1 Let $X_{\sigma}=\operatorname{Spec}\mathbb{C}[\sigma^{\vee}\cap M]$ be the affine toric variety corresponding to a strongly convex rational polyhedral cone $\sigma$. For an $r$-dimensional face $\tau\subset\sigma$, we get a torus invariant subvariety $S_{\tau}=\operatorname{Spec}\mathbb{C}[\sigma^{\vee}\cap\tau^{\perp}\cap M]\subset X_{\sigma}$ of codimension $r$. This is the affine toric variety corresponding to the cone $\bar{\sigma}_{\tau}$, where the lattice and the dual lattice are given by
    $N_{\tau}:=\frac{N}{N\cap\langle\tau\rangle},\qquad M_{\tau}:=M\cap\tau^{\perp}.$
    We denote by $O_{\tau}=\operatorname{Spec}\mathbb{C}[M_{\tau}]$ the torus orbit corresponding to $\tau$. Also, $U_{\tau}=\operatorname{Spec}\mathbb{C}[\tau^{\vee}\cap M]$ is an open subset of $X_{\sigma}$ and we have the diagram of torus equivariant morphisms
[…]
    After fixing a non-canonical splitting $N=N_{\tau}\oplus(N\cap\langle\tau\rangle)$ and the corresponding splitting $M=M_{\tau}\oplus M^{\prime}$, we can identify $U_{\tau}\to O_{\tau}$ as the projection $U_{\tau}=V_{\tau}\times O_{\tau}\to O_{\tau}$, where $V_{\tau}$ is the full-dimensional toric variety $\operatorname{Spec}\mathbb{C}[\tau^{\vee}\cap M^{\prime}]$.
    For two faces $\mu\subset\tau$, we denote by $V_{\mu,\tau}$ the full-dimensional affine toric variety corresponding to the cone $\bar{\tau}_{\mu}\subset\langle\tau\rangle/\langle\mu\rangle$. We have an analogous diagram
[…]
    Note that
    $\Sigma_{S_{\mu}}$ =$=\{\bar{\tau}_{\mu}\colon\tau\in\Sigma_{X},\mu\subset\tau\},$
    $\Sigma_{V_{\tau}}$ =$=\{\mu\in\Sigma_{X}\colon\mu\subset\tau\},$
    $\Sigma_{V_{\mu,\tau}}$ =$=\{\bar{\nu}_{\mu}\colon\nu\in\Sigma_{X},\mu\subset\nu\subset\tau\}$
@@
    We say that a toric variety $X$ is simplicial if all the cones $\sigma$ in the fan are simplicial. If $X$ is simplicial, then $X$ has quotient singularities [CoxLittleSchenck-ToricVar, Theorem 11.4.8]. By Lemma 2.3, in this case we have a canonical isomorphism
    $\operatorname{gr}_{k}\operatorname{DR}_{X}\operatorname{IC}_{X}^{H}\simeq\Omega_{X}^{[-k]}[n+k].$
@@
    [oda-ConvexBodies, Lemma 3.5] (see also [Ishida]) There is a natural resolution of the sheaf of reflexive differentials $\Omega_{Y}^{[p]}:=(\Omega_{Y}^{p})^{\vee\vee}$ of a simplicial toric variety $Y$, called the Ishida complex, that we now recall. Let $\tau\in\Sigma_{Y}$ be an $r$-dimensional cone. Since $Y$ is simplicial, $\tau$ is generated by $r$ rays $\rho_{1},\ldots,\rho_{r}$. We set
    $V_{\tau}^{p}:=\left(\bigwedge^{p-r}\tau^{\perp}\right)\otimes\frac{M_{\mathbb{C}}}{\rho_{1}^{\perp}}\otimes\ldots\otimes\frac{M_{\mathbb{C}}}{\rho_{r}^{\perp}},$
    where $M_{\mathbb{C}}:=M\otimes_{\mathbb{Z}}\mathbb{C}$. For instance, $V_{0}^{p}=\bigwedge^{p}M_{\mathbb{C}}$. Then we have an exact complex
    (2) $0\to\Omega_{Y}^{[p]}\to V_{0}^{p}\otimes\mathcal{O}_{Y}\to\bigoplus_{\begin{subarray}{c}\tau\in\Sigma_{Y}\\ \dim\tau=1\end{subarray}}V_{\tau}^{p}\otimes\mathcal{O}_{S_{\tau}}\to\bigoplus_{\begin{subarray}{c}\tau\in\Sigma_{Y}\\ \dim\tau=2\end{subarray}}V_{\tau}^{p}\otimes\mathcal{O}_{S_{\tau}}\to\ldots\to\bigoplus_{\begin{subarray}{c}\tau\in\Sigma_{Y}\\ \dim\tau=p\end{subarray}}V_{\tau}^{p}\otimes\mathcal{O}_{S_{\tau}}\to 0$
@@
where $F^{(i)}$ is a facet of $F^{(i+1)}$ for $10\leq i\leq n-$ 1. Observe that $F^{(0)}$ is a vertex of $P$. Given this sequence, the set of vertices of the corresponding simplex consists of
@@ 2.3. Polytopes and Shellability
    For $1<j\leq s$,
    $F_{j}\cap\left(\bigcup_{i=1}^{j-1}F_{i}\right)=G_{1}\cup G_{2}\cup\ldots\cup G_{r}$
    for some shelling $G_{1},G_{2}\ldots,G_{r},\ldots,G_{t}$ of $\mathcal{C}(\partial F_{j})$
@@
Let $F_{1},\dots,F_{r}$ be a shelling of a simplicial polytopal complex. Define $F_{1}$ to be of type 00. For $j\geq 2$, we have that $F_{j}\cap(\bigcup_{i=1}^{j-1}F_{i})$ is a pure $(d-1)$-dimensional complex. Define $F_{j}$ to be of type $l$, where $l$ is the number of facets in the pure complex $F_{j}\cap(\bigcup_{i=1}^{j-1}F_{i})$.
@@
{Faces of Fj containing [$\left\{\text{Faces of $F_{j}$ containing $[v_{j_{1}},\ldots,v_{j_{l}}]$}\right\}$ =$=\left\{\text{Faces of $F_{j}$ not contained in any $G_{k}$ $\forall k=1,\dots,l$}\right\}$
=$=\left\{\text{Faces of $F_{j}$ not contained in $\bigcup_{k=1}^{l}G_{k}$}\right\}$
=$=\left\{\text{Faces of $F_{j}$ not contained in $F_{j}\cap\left(\bigcup_{i=1}^{j-1}F_{i}\right)$}\right\}.$
@@
where $F^{(i)}$ is a facet of $F^{(i+1)}$ for $10\leq i\leq n-$ 1. Observe that $F^{(0)}$ is a vertex of $P$. From now on, for a maximal simplex $\Delta$ in $\mathcal{C}_{Y}$ corresponding to the chain of faces $P=F^{(n)},\ldots,F^{(0)}$, we use the notation
@@
Similarly, given an $n-1$ dimensional face $F^{(n-1)}$ of $P$, let us now define an order $\prec_{F^{(n)},F^{(n-1)}}$ on the set of facets of $F^{(n-1)}$. Observe that
@@
Note that the facets $[v_{(n)},\ldots,\widehat{v_{(i_{t})}},\ldots,v_{(0)}]$ are the facets of $\Delta$ which are contained in $\bigcup_{\Delta^{\prime}\prec\Delta}\Delta^{\prime}$. We claim that any face of $\Delta$ containing $[v_{(i_{1})},\ldots,v_{(i_{t})}]$ is not a face of $\Delta^{\prime}$ with $\Delta^{\prime}\prec\Delta$. Provided that, it is straightforward to see that
@@
We are left with proving that any face of $\Delta$ containing $[v_{(i_{1})},\ldots,v_{(i_{t})}]$ is not a face of $\Delta^{\prime}$ with $\Delta^{\prime}\prec\Delta$, that is, we want to prove that the set $S$ of simplices $\Delta^{\prime}=(G^{(n)}\supset\ldots\supset G^{(0)})$ satisfying the following properties:
@@
First, we suppose that $i>i_{l}$. We have $G^{(i)}\prec_{F^{(n)},\ldots,F^{(i+1)}}F^{(i)}$. Also, we have $F^{(i_{l})}=G^{(i_{l})}$ which implies that $F^{(i)}\cap G^{(i)}$ contains $F^{(i_{l})}$. This implies that there exists a facet $\widetilde{G}^{(i)}\prec_{F^{(n)},\ldots,F^{(i+1)}}F^{(i)}$ of $F^{(i+1)}$ such that $\widetilde{G}^{(i-1)}:=\widetilde{G}^{(i)}\cap F^{(i)}$ is a $(i-1)$-dimensional face and $F^{(i_{l})}$ is a face of $\widetilde{G}^{(i-1)}$. We replace $\Delta^{\prime}$ with
@@ 2.4. The Decomposition Theorem
We finally upgrade the decomposition in (3) at the level of Hodge modules. Given a Hodge module $\mathcal{M}$ of weight $w$ with filtration $F_{\bullet}$ and underlying perverse sheaf $K$, the Tate twist of $M$ by an integer $k$ is a Hodge module, denoted $\mathcal{M}(k)$, of weight $w-2k$ which has the same $\mathcal{D}$-module structure as $\mathcal{M}$ but with the filtration $F_{\bullet-k}$ and the perverse sheaf $K\otimes_{\mathbb{Q}}(2\pi i)^{k}\mathbb{Q}$.
@@
[…]
@@ 3. Higher direct images of Kähler differentials for toric morphisms
(8) $0\to\Omega_{X}^{[p]}\to V_{0}^{p}\otimes\mathcal{O}_{X}\to\bigoplus_{\begin{subarray}{c}\nu\in\Sigma_{Y}\\ \dim\nu=1\end{subarray}}V_{\nu}^{p}\otimes\mathcal{O}_{S_{\pi_{*}(\nu)}}\to\bigoplus_{\begin{subarray}{c}\nu\in\Sigma_{Y}\\ \dim\nu=2\end{subarray}}V_{\nu}^{p}\otimes\mathcal{O}_{S_{\pi_{*}(\nu)}}\to\ldots\to\bigoplus_{\begin{subarray}{c}\nu\in\Sigma_{Y}\\ \dim\nu=p\end{subarray}}V_{\nu}^{p}\otimes\mathcal{O}_{S_{\pi_{*}(\nu)}}\to 0$
@@
$\left(\Omega_{X}^{[p]}\right)_{u}=\{\alpha\in\Omega_{X}^{[p]}:g^{*}\alpha=u(g)\cdot\alpha\text{ for all }g\in\operatorname{Spec}\mathbb{C}[M]\}\$
@@
with $V^{p}_{0}$ in cohomological degree 00.
@@
(10) $0\to V_{0}^{p}\to\bigoplus_{\begin{subarray}{c}\gamma\in\mathcal{C}_{Y}\\ \dim\gamma=0\end{subarray}}V_{\gamma}^{p}\to\bigoplus_{\begin{subarray}{c}\gamma\in\mathcal{C}_{Y}\\ \dim\gamma=1\end{subarray}}V_{\gamma}^{p}\to\ldots\to\bigoplus_{\begin{subarray}{c}\gamma\in\mathcal{C}_{Y}\\ \dim\gamma=p-1\end{subarray}}V_{\gamma}^{p}\to 0$
@@
with $V_{0}^{p}$ in cohomological degree 00, by putting a filtration on $A^{\bullet}$ from the shelling and computing the associated spectral sequence. We define $F^{i}A^{\bullet}$ by
@@
(12) $0\to V_{0}^{p}\to\bigoplus_{\begin{subarray}{c}\nu\in\Sigma_{Y}\\ \nu\subset\tau,\dim\nu=1\end{subarray}}V_{\nu}^{p}\to\bigoplus_{\begin{subarray}{c}\nu\in\Sigma_{Y}\\ \nu\subset\tau,\dim\nu=2\end{subarray}}V_{\nu}^{p}\to\ldots\to\bigoplus_{\begin{subarray}{c}\nu\in\Sigma_{Y}\\ \nu\subset\tau,\dim\nu=p\end{subarray}}V_{\nu}^{p}\to 0$
@@
is concentrated in degree $i$ and is equal to $(R^{i}\pi_{*}\Omega_{Y}^{[p]})_{u}$. Furthermore, the number of summands of the $j$-th term is equal to $\sum_{\mu\subset\tau,\mu\in\Sigma_{X}}d_{j}(\mu)$ and the dimension of each summand of the $j$-th term is equal to (${d_{\tau}-j\choose i-j}$. Hence we have
@@
=($={n-d_{\tau}\choose p-i}\cdot\left((-1)^{i}{d_{\tau}\choose i}\sum_{\begin{subarray}{c}\mu\subset\tau\\ \mu\in\Sigma_{X}\end{subarray}}d_{0}(\mu)+(-1)^{i-1}{d_{\tau}-1\choose i-1}\sum_{\begin{subarray}{c}\mu\subset\tau\\ \mu\in\Sigma_{X}\end{subarray}}d_{1}(\mu)+\cdots+{d_{\tau}-i\choose 0}\sum_{\begin{subarray}{c}\mu\subset\tau\\ \mu\in\Sigma_{X}\end{subarray}}d_{i}(\mu)\right)$
=($={n-d_{\tau}\choose p-i}\sum_{j=0}^{i}\sum_{\begin{subarray}{c}\mu\subset\tau\\ \mu\in\Sigma_{X}\end{subarray}}(-1)^{i-j}{d_{\tau}-j\choose i-j}d_{j}(\mu).$
@@
(where we set $\qquad(\text{where we set }c=2k+l)$
@@
(where we set $\qquad(\text{where we set }k^{\prime}=n+c-j-k)$
@@
(where we set $\qquad(\text{where we set }c^{\prime}=-n-c)$

### v4

@@ 1. Introduction
Remark 1.2.
After the submission of our paper, we were informed by the referee of a possible alternative way to prove Theorem 1.1. We properly spell this out in Appendix B.
Although the approach in the appendix is a simpler way to prove Theorem 1.1, we believe our approach yields several additional results along the way (e.g. developing a systematic way to calculate higher direct images of reflexive differential forms for toric morphisms, using shellability in these computations), which are interesting in their own right. In fact, we build on many of these ideas to provide a systematic understanding of the trivial Hodge module $\mathbb{Q}_{X}^{H}[n]$ on toric varieties, the local cohomology modules of toric varieties, and the Hodge structure on the singular cohomology of toric varieties in follow-up works [LCDTV2, LCDTV1].
@@ 2.1. Hodge modules
which sits in cohomological degrees $\dim X,\ldots,0$. Moreover, $F_{\bullet}$ induces a natural filtration on $\operatorname{DR}_{X}(\mathcal{M})$ and the graded pieces $\operatorname{gr}_{k}\operatorname{DR}_{X}\mathcal{M}$ are given by
@@
which also sits in cohomological degrees $\dim X,\ldots,0$. The maps in this complex are $\mathcal{O}_{X}$-linear and we thus view $\operatorname{gr}_{k}\operatorname{DR}_{X}\mathcal{M}$ as an object in $\mathcal{D}^{b}_{\rm coh}(X)$, the derived category of coherent sheaves of $X$. We mention that even if $X$ is singular, one can define the categories $\operatorname{HM}(X,w)$ and $\operatorname{MHM}(X)$ by embedding into a smooth variety $X\hookrightarrow Y$ and considering the objects in $\operatorname{HM}(Y,w)$ and $\operatorname{MHM}(Y)$, respectively, which are supported on $X$. The categories $\operatorname{HM}(X,w)$ and $\operatorname{MHM}(X)$ do not depend on the choice of the embedding. If $X$ cannot be embedded in a smooth variety, we locally embed each open set and impose suitable compatibility conditions on the intersections. The de Rham complex and the graded pieces $\operatorname{gr}_{k}\operatorname{DR}_{X}\mathcal{M}$ are defined as above by locally embedding $X$ into a smooth variety. The graded pieces of the de Rham complex $\operatorname{gr}_{k}\operatorname{DR}_{X}\mathcal{M}$ also do not depend on the choice of the embedding as objects in $\mathcal{D}_{\operatorname{coh}}^{b}(X)$.
@@ 2.2. Toric varieties
    [Fulton-ToricVar]*Section 3.1 Let $X_{\sigma}=\operatorname{Spec}\mathbb{C}[\sigma^{\vee}\cap M]$ be the affine toric variety corresponding to a strongly convex rational polyhedral cone $\sigma$. For an $r$-dimensional face $\tau\subset\sigma$, we get a torus invariant subvariety $S_{\tau}=\operatorname{Spec}\mathbb{C}[\sigma^{\vee}\cap\tau^{\perp}\cap M]\subset X_{\sigma}$ of codimension $r$. This is the affine toric variety corresponding to the cone $\bar{\sigma}_{\tau}$, where the lattice and the dual lattice are given by N_τ := NN ∩⟨τ⟩,  M_τ: = M ∩τ^⟂. We denote by $O_{\tau}=\operatorname{Spec}\mathbb{C}[M_{\tau}]$ the torus orbit corresponding to $\tau$. Also, $U_{\tau}=\operatorname{Spec}\mathbb{C}[\tau^{\vee}\cap M]$ is an open subset of $X_{\sigma}$ and we have the diagram of torus equivariant morphisms $U_{\tau}$$X_{\sigma}$$O_{\tau}$$S_{\tau}.$} After fixing a non-canonical splitting $N=N_{\tau}\oplus(N\cap\langle\tau\rangle)$ and the corresponding splitting $M=M_{\tau}\oplus M^{\prime}$, we can identify $U_{\tau}\to O_{\tau}$ as the projection $U_{\tau}=V_{\tau}\times O_{\tau}\to O_{\tau}$, where $V_{\tau}$ is the full-dimensional toric variety $\operatorname{Spec}\mathbb{C}[\tau^{\vee}\cap M^{\prime}]$.
    For two faces $\mu\subset\tau$, we denote by $V_{\mu,\tau}$ the full-dimensional affine toric variety corresponding to the cone $\bar{\tau}_{\mu}\subset\langle\tau\rangle/\langle\mu\rangle$. We have an analogous diagram $S_{\mu}$$U_{\bar{\tau}_{\mu}}\simeq V_{\mu,\tau}\times O_{\tau}$$S_{\tau}$$O_{\tau}.$} Note that
    $\Sigma_{S_{\mu}}$ $=\{\bar{\tau}_{\mu}\colon\tau\in\Sigma_{X},\mu\subset\tau\},$
    $\Sigma_{V_{\tau}}$ $=\{\mu\in\Sigma_{X}\colon\mu\subset\tau\},$
    $\Sigma_{V_{\mu,\tau}}$ $=\{\bar{\nu}_{\mu}\colon\nu\in\Sigma_{X},\mu\subset\nu\subset\tau\}$
@@
    We say that a toric variety $X$ is simplicial if all the cones $\sigma$ in the fan are simplicial. If $X$ is simplicial, then $X$ has quotient singularities [CoxLittleSchenck-ToricVar, Theorem 11.4.8]. By Lemma 2.3, in this case we have a canonical isomorphism gr_kDR_X IC_X^H ≃Ω_X^[-k][n+k].
@@
    [oda-ConvexBodies, Lemma 3.5] (see also [Ishida]) There is a natural resolution of the sheaf of reflexive differentials $\Omega_{Y}^{[p]}:=(\Omega_{Y}^{p})^{\vee\vee}$ of a simplicial toric variety $Y$, called the Ishida complex, that we now recall. Let $\tau\in\Sigma_{Y}$ be an $r$-dimensional cone. Since $Y$ is simplicial, $\tau$ is generated by $r$ rays $\rho_{1},\ldots,\rho_{r}$. We set V_τ^p := ( ⋀^p-r τ^⟂ ) ⊗MCρ1 ⊗…⊗MCρr, where $M_{\mathbb{C}}:=M\otimes_{\mathbb{Z}}\mathbb{C}$. For instance, $V_{0}^{p}=\bigwedge^{p}M_{\mathbb{C}}$. Then we have an exact complex
    (2) $0\to\Omega_{Y}^{[p]}\to V_{0}^{p}\otimes\mathcal{O}_{Y}\to\bigoplus_{\begin{subarray}{c}\tau\in\Sigma_{Y}\\ \dim\tau=1\end{subarray}}V_{\tau}^{p}\otimes\mathcal{O}_{S_{\tau}}\to\bigoplus_{\begin{subarray}{c}\tau\in\Sigma_{Y}\\ \dim\tau=2\end{subarray}}V_{\tau}^{p}\otimes\mathcal{O}_{S_{\tau}}\to\ldots\to\bigoplus_{\begin{subarray}{c}\tau\in\Sigma_{Y}\\ \dim\tau=p\end{subarray}}V_{\tau}^{p}\otimes\mathcal{O}_{S_{\tau}}\to 0.$
@@
where $F^{(i)}$ is a facet of $F^{(i+1)}$ for $0\leq i\leq n-1$. Observe that $F^{(0)}$ is a vertex of $P$. Given this sequence, the set of vertices of the corresponding simplex consists of
@@ 2.3. Polytopes and Shellability
    For $1<j\leq s$, F_j ∩( ⋃_i=1^j-1 F_i ) = G_1 ∪G_2 ∪…∪G_r for some shelling $G_{1},G_{2}\ldots,G_{r},\ldots,G_{t}$ of $\mathcal{C}(\partial F_{j})$
@@
Let $F_{1},\dots,F_{r}$ be a shelling of a simplicial polytopal complex. Define $F_{1}$ to be of type 0. For $j\geq 2$, we have that $F_{j}\cap(\bigcup_{i=1}^{j-1}F_{i})$ is a pure (d1)(d-1)-dimensional complex. Define $F_{j}$ to be of type $l$, where $l$ is the number of facets in the pure complex $F_{j}\cap(\bigcup_{i=1}^{j-1}F_{i})$.
@@
[…]
$=\left\{\text{Faces of $F_{j}$ not contained in $\bigcup_{k=1}^{l}G_{k}$}\right\}$
$=\left\{\text{Faces of $F_{j}$ not contained in $F_{j}\cap\left(\bigcup_{i=1}^{j-1}F_{i}\right)$}\right\}.$
@@
where $F^{(i)}$ is a facet of $F^{(i+1)}$ for $0\leq i\leq n-1$. Observe that $F^{(0)}$ is a vertex of $P$. From now on, for a maximal simplex $\Delta$ in $\mathcal{C}_{Y}$ corresponding to the chain of faces $P=F^{(n)},\ldots,F^{(0)}$, we use the notation
@@
Similarly, given an n1n-1 dimensional face $F^{(n-1)}$ of $P$, let us now define an order $\prec_{F^{(n)},F^{(n-1)}}$ on the set of facets of $F^{(n-1)}$. Observe that
@@
Note that the facets $v_{(n)},\ldots,\widehat{v_{(i_{t})}},\ldots,v_{(0)}]$ are the facets of $\Delta$ which are contained in $\bigcup_{\Delta^{\prime}\prec\Delta}\Delta^{\prime}$. We claim that any face of $\Delta$ containing $v_{(i_{1})},\ldots,v_{(i_{t})}]$ is not a face of $\Delta^{\prime}$ with $\Delta^{\prime}\prec\Delta$. Provided that, it is straightforward to see that
@@
We are left with proving that any face of $\Delta$ containing $v_{(i_{1})},\ldots,v_{(i_{t})}]$ is not a face of $\Delta^{\prime}$ with $\Delta^{\prime}\prec\Delta$, that is, we want to prove that the set $S$ of simplices $\Delta^{\prime}=(G^{(n)}\supset\ldots\supset G^{(0)})$ satisfying the following properties:
@@
First, we suppose that $i>i_{l}$. We have $G^{(i)}\prec_{F^{(n)},\ldots,F^{(i+1)}}F^{(i)}$. Also, we have $F^{(i_{l})}=G^{(i_{l})}$ which implies that $F^{(i)}\cap G^{(i)}$ contains $F^{(i_{l})}$. This implies that there exists a facet $\widetilde{G}^{(i)}\prec_{F^{(n)},\ldots,F^{(i+1)}}F^{(i)}$ of $F^{(i+1)}$ such that $\widetilde{G}^{(i-1)}:=\widetilde{G}^{(i)}\cap F^{(i)}$ is a (i1)(i-1)-dimensional face and $F^{(i_{l})}$ is a face of $\widetilde{G}^{(i-1)}$. We replace $\Delta^{\prime}$ with
@@ 2.4. The Decomposition Theorem
We finally upgrade the decomposition in (3) at the level of Hodge modules. Given a Hodge module $\mathcal{M}$ of weight $w$ with filtration $F_{\bullet}$ and underlying perverse sheaf $K$, the Tate twist of $M$ by an integer $k$ is a Hodge module, denoted $\mathcal{M}(k)$, of weight w2kw-2k which has the same $\mathcal{D}$-module structure as $\mathcal{M}$ but with the filtration $F_{\bullet-k}$ and the perverse sheaf $K\otimes_{\mathbb{Q}}(2\pi i)^{k}\mathbb{Q}$.
@@
$Y_{\mu_{0}}$$W_{\mu_{0}}$$U_{\mu_{0}}$$V_{\mu_{0}}$$\scriptstyle{\operatorname{pr}}$$\scriptstyle{\pi}$$\scriptstyle{p}$$\scriptstyle{\operatorname{pr}}$
@@ 3. Higher direct images of Kähler differentials for toric morphisms
(8) $0\to\Omega_{X}^{[p]}\to V_{0}^{p}\otimes\mathcal{O}_{X}\to\bigoplus_{\begin{subarray}{c}\nu\in\Sigma_{Y}\\ \dim\nu=1\end{subarray}}V_{\nu}^{p}\otimes\mathcal{O}_{S_{\pi_{*}(\nu)}}\to\bigoplus_{\begin{subarray}{c}\nu\in\Sigma_{Y}\\ \dim\nu=2\end{subarray}}V_{\nu}^{p}\otimes\mathcal{O}_{S_{\pi_{*}(\nu)}}\to\ldots\to\bigoplus_{\begin{subarray}{c}\nu\in\Sigma_{Y}\\ \dim\nu=p\end{subarray}}V_{\nu}^{p}\otimes\mathcal{O}_{S_{\pi_{*}(\nu)}}\to 0.$
@@
$\left(\Omega_{X}^{[p]}\right)_{u}=\{\alpha\in\Omega_{X}^{[p]}:g^{*}\alpha=u(g)\cdot\alpha\text{ for all }g\in\operatorname{Spec}\mathbb{C}[M]\}\\$
@@
with $V^{p}_{0}$ in cohomological degree 0.
@@
(10) $0\to V_{0}^{p}\to\bigoplus_{\begin{subarray}{c}\gamma\in\mathcal{C}_{Y}\\ \dim\gamma=0\end{subarray}}V_{\gamma}^{p}\to\bigoplus_{\begin{subarray}{c}\gamma\in\mathcal{C}_{Y}\\ \dim\gamma=1\end{subarray}}V_{\gamma}^{p}\to\ldots\to\bigoplus_{\begin{subarray}{c}\gamma\in\mathcal{C}_{Y}\\ \dim\gamma=p-1\end{subarray}}V_{\gamma}^{p}\to 0.$
@@
with $V_{0}^{p}$ in cohomological degree 0, by putting a filtration on $A^{\bullet}$ from the shelling and computing the associated spectral sequence. We define $F^{i}A^{\bullet}$ by
@@
(12) $0\to V_{0}^{p}\to\bigoplus_{\begin{subarray}{c}\nu\in\Sigma_{Y}\\ \nu\subset\tau,\dim\nu=1\end{subarray}}V_{\nu}^{p}\to\bigoplus_{\begin{subarray}{c}\nu\in\Sigma_{Y}\\ \nu\subset\tau,\dim\nu=2\end{subarray}}V_{\nu}^{p}\to\ldots\to\bigoplus_{\begin{subarray}{c}\nu\in\Sigma_{Y}\\ \nu\subset\tau,\dim\nu=p\end{subarray}}V_{\nu}^{p}\to 0.$
@@
is concentrated in degree $i$ and is equal to $(R^{i}\pi_{*}\Omega_{Y}^{[p]})_{u}$. Furthermore, the number of summands of the $j$-th term is equal to $\sum_{\mu\subset\tau,\mu\in\Sigma_{X}}d_{j}(\mu)$ and the dimension of each summand of the $j$-th term is equal to $d_{\tau}-j\choose i-j$}. Hence we have
@@
$={n-d_{\tau}\choose p-i}\cdot\left((-1)^{i}{d_{\tau}\choose i}\sum_{\begin{subarray}{c}\mu\subset\tau\\ \mu\in\Sigma_{X}\end{subarray}}d_{0}(\mu)+(-1)^{i-1}{d_{\tau}-1\choose i-1}\sum_{\begin{subarray}{c}\mu\subset\tau\\ \mu\in\Sigma_{X}\end{subarray}}d_{1}(\mu)+\cdots+{d_{\tau}-i\choose 0}\sum_{\begin{subarray}{c}\mu\subset\tau\\ \mu\in\Sigma_{X}\end{subarray}}d_{i}(\mu)\right)$
$={n-d_{\tau}\choose p-i}\sum_{j=0}^{i}\sum_{\begin{subarray}{c}\mu\subset\tau\\ \mu\in\Sigma_{X}\end{subarray}}(-1)^{i-j}{d_{\tau}-j\choose i-j}d_{j}(\mu).$
@@
$\qquad(\text{where we set }c=2k+l)$
@@
$\qquad(\text{where we set }k^{\prime}=n+c-j-k)$
@@
$\qquad(\text{where we set }c^{\prime}=-n-c)$